# Shell quenching in nuclear charge radii based on Monte Carlo dropout Bayesian neural network

### v1

@@ III Results and Discussion
For the convenience of discussion, we use “$P^{*}$” to denote the model combining NP formula and the MC-dropout BNN with input neurons $z=\{Z,$ $A$, $I^{2}$, $\delta$, $P^{*}$, $\nu_{n}$, $\nu_{p}$, $\beta_{20}$, $\beta_{40}$, $LI\}$. We also give the results obtained by the MC-dropout BNN with input neurons $z=\{Z,$ $A$, $I^{2}$, $\delta$, $P$, $\nu_{n}$, $\nu_{p}$, $\beta_{20}$, $\beta_{40}$, $LI\}$, denoted by “$P$” for comparison. Here, we should be mentioned that the difference between $P^{*}$ and $P$ just distinguishes the local variations in charge radii along $Z=28$, $50$, $82$ isotopic families. In the training set, we use the $815$ experimental data given in Ref. ANGELI201369 . The more recent experimental data PhysRevLett.126.032502 ; LI2021101440 , containing 118 data for nuclei with $Z\geq 20$ and $A\geq 40$, are used as the validation set to test the predictive power of our model. In contrast to Refs. PhysRevC.105.014308 ; DONG2023137726 , the training data sets are reduced owing to the whole data of charge radii along nickel isotopic chain are transformed into the validation data set.
@@
To further quantify the predictive power, the rms deviation between the results obtained by the MC-dropout BNN model and the corresponding experimental data are calculated as follows,
$\sigma^{(\mathrm{T},\mathrm{V},\mathrm{TV})}=\sqrt{\frac{1}{N_{(\mathrm{T},\mathrm{V},\mathrm{TV})}}\sum_{i=1}^{N_{(\mathrm{T},\mathrm{V},\mathrm{TV})}}\left(R_{\mathrm{ch},i}^{\mathrm{pred}.}-R_{\mathrm{ch},i}^{\mathrm{exp}.}\right)^{2}}.$ (11)
@@
where $N_{\mathrm{T}}$, $N_{\mathrm{V}}$, and $N_{\mathrm{TV}}$ are the numbers of charge radii contained in the training (T), validation (V), and entire (TV) data sets, and the subscript $i$ represents the $i$th nucleus in the given data sets.
@@
As shown in Ref. PhysRevC.105.014308 , the predictive accuracy in describing the rms charge radii has been achieved to be $\sigma^{\mathrm{T}}=0.0143$ fm, and $\sigma^{\mathrm{TV}}=0.0149$ fm. Then the revised version has been proposed to further reduce the quantitative uncertainty DONG2023137726 , which leads to the $\sigma^{\mathrm{T}}=0.0140$ fm and $\sigma^{\mathrm{TV}}=0.0139$ fm. For these studies, the Casten factor associated to the valence neutrons and protons is regarded as input parameter. However, the quadrupole deformation has not been taken into account in the validated protocol. As demonstrated in Ref. PhysRevC.88.011301 , the rms deviation in charge radii can be reduced if the shape deformation is incorporated into the calibrated protocol. In recent study, the input sets $\{Z,A,\delta,\nu_{p},\nu_{n},\beta_{20}\}$ have been used in the training and validation data set, the rms deviation has been fallen into $\sigma^{\mathrm{T}}=0.0106$ fm and $\sigma^{\mathrm{TV}}=0.0130$ fm Zhang10662945 .
@@
In Table 1, the $\sigma^{\mathrm{T}}$, $\sigma^{\mathrm{V}}$, and $\sigma^{\mathrm{TV}}$ values derived from the Eq. (11) are shown, respectively. In our calculations, the valence neutrons, valence protons, Casten factor, and the shape-phase deformation in the limited region are simultaneously taken into account during the simulated process. Meanwhile, the higher deformation parameter $\beta_{40}$ is also included. The deduced value of $\sigma^{\mathrm{TV}}$ deriving from the $P$ model is equivalent to Ref. Zhang10662945 , but less than the results shown in Refs. PhysRevC.105.014308 ; DONG2023137726 . As mentioned above, the over-fitting puzzle can be encountered in the traditional Casten factor $P$ when meets the isotopic chains with proton numbers $Z=28$, $50$, $82$. The $\sigma^{\mathrm{TV}}$ value obtained by $P^{*}$ model can be further reduced to be $0.0094$ fm. This is accord with Ref. DONG2023137726 where the modified rms deviation has been used to measure the quantified uncertainty. In a word, the rms deviations in the training and validation data sets are also further reduced with respect to those investigations in the existing literature. It is therefore reasonable to demonstrate the predictive power of the $P^{*}$ model by modifying the traditional Casten factor $P$.
Significant improvement has been achieved by the $P^{*}$ model in both the interpolation and the extrapolation predictions of nuclear charge radii in comparison with the $P$ model. To visualize the details about the training and validation of the MC-dropout BNN approach, as shown in Fig. 2, the difference between the calibrated rms charge radii with $P^{*}$ model and the corresponding experimental data are depicted in nuclear chart. From this figure, it can be found that most of the experimental data can be reproduced well. In some specific regions, the predicted values are heavily deviated from the experimental data. As is well known, the inverse OES or the weakened OES can be evidently observed around Eu ($Z=63$Alkhazov1990 and At($Z=85$PhysRevC.99.054317 isotopes. In our calibrated protocol, the inverse odd-even staggering behavior has not been considered properly. This seems to lead the larger deviations between the predicted rms charge radii and the corresponding experimental data. Meanwhile, the larger deviations can also be encountered along manganese isotopes. This means that more underlying mechanisms should be covered through the input structures.
@@
Figure 2: (Color online)Nuclear chart displaying the absolute differences (in units of fm) between the MC-Dropout BNN predicted rms charge radii and the experimental data in the entire set. The results are obtained from the training and validation procedures, respectively.
@@
Figure 3: (Color online)Root-mean-square charge radii of calcium (a), nickel (b), tin (c), and lead (d) isotopes simulated by the $P^{*}$ and $P$ models, in comparison with the experimental data ANGELI201369 ; LI2021101440 . The results obtained by the np formula ($R_{\mathrm{np}}$) are also drawn for making a guide eye. The data covered by gray area are not contained in the training set, namely the pure predictions. The neutron magic numbers $N=20$, 28, 50, 82, and 126 are emphasized by the vertical lines.
@@
Shell closure effects are naturally observed in charge radii throughout the whole nuclear chart ANGELI201369 ; LI2021101440 ; PhysRevLett.128.022502 ; PhysRevLett.129.132501 ; PhysRevLett.131.102501 . This attributes to the rather small isospin dependence of the spin-orbit interactions PhysRevLett.74.3744 . The relativistic mean field theory Geng:2003pk ; PhysRevC.102.024307 ; PhysRevC.109.064302 and the Fayans energy density functional model PhysRevC.95.064328 ; PhysRevLett.122.192502 can reproduce the shell quenching phenomena in nuclear charge radii well. However, as demonstrated in Ref. PhysRevLett.128.022502 , the Skyrme density functional theory cannot reproduce the shell closure effects. This provides an potential signature to identify the predictive power of nuclear models in describing the local variations of charge radii. In Fig. 3, the rms charge radii of calcium (a), nickel (b), tin (c), and lead (d) isotopes are depicted with both of the $P^{*}$ and $P$ models, respectively. As mentioned above, MC-dropout BNN model can provide the quantified uncertainty in the validated process. Therefore the corresponding error bars are also shown in Fig. 3. For calcium isotopes, $P^{*}$ and $P$ models give the similar trend of changes of charge radii. The slight deviations can be found at the neutron-deficient 34Ca and neutron-rich 54,55Ca isotopes. Meanwhile, charge radii of 39,43Ca isotopes can be reproduced well by the $P^{*}$ model in comparing with the $P$ model. Likewise, the systematical evolution of charge radii along tin isotopes can also be reproduced well by both of the $P^{*}$ and $P$ models as shown in Fig. 3 (c).
@@
In Fig. 3 (b), it is obviously found that the $N=28$ shell closure effect cannot be reproduced well by the $P$ model. Besides, the charge radii of 54-56,58Ni isotopes are heavily overestimated. The same scenarios can also be found in the 65-68Ni isotopes, but with slight deviation. Meanwhile, charge radius of 70Ni obtained by the $P$ model is slightly underestimated with respect to $P^{*}$ model. Toward proton-rich regions, the pronounced deviations can be obviously exhibited between the predictive results obtained by the $P^{*}$ and $P$ models. These larger deviations can also be encountered in the neutron-rich lead isotopes as shown in Fig. 3 (d). Here, it should be mentioned that the rapid increase of charge radii can be reproduced well by the $P^{*}$ model, but the $P$ model fails to follow this trend beyond the neutron number $N=126$. All of these results suggest that $P^{*}$ model can provide a powerful prediction of nuclear charge radii. Therefore, the extrapolated simulation with high accuracy is necessary. In Table. 2, charge radii of calcium, nickel, tin, and lead isotopes are shown explicitly. These results are encouraged to provide an theoretical guidance for the experimental researches.
@@
The input parameters which include the modified Casten factor $P^{*}$ can describe the systematical evolution of nuclear charge radii well. This can be easily found from the rms deviation of the training and validation data sets. Furthermore, as shown in Fig. 3, the shrunk phenomena in charge radii can be reproduced well with the $P^{*}$ model. Actually, Casten factor represents the neutron-proton correlation in describing the trend of changes of nuclear charge radii. As demonstrated in Refs. PhysRevC.53.1599 ; MILLER2019360 ; COSYN2021136526 , the neutron-proton correlation has an influence on determining the nuclear charge radii. This can also be found from Refs. PhysRevC.102.024307 ; PhysRevC.109.064302 where the neutron-proton correlation derived from the difference of the fractions of Cooper pairs between neutrons and protons can reproduce the shell quenching in charge radii as well. All of these suggest that more appropriate isospin interactions should be covered in describing the nuclear charge radii. Especially, the difference of charge radii can be used to ascertain the nuclear symmetry energy through the mirror-pair nuclei. Thus more available charge radii data are required in theoretical study.
@@ IV Summary and conclusions
More input parameter sets physically motivated, such as proton number, mass number, isospin-asymmetry degree, pairing effect, shell closure effect, valence neutrons and protons, Casten factor $P$ or $P^{*}$, local shape-staggering phenomena, quadrupole deformation $\beta_{20}$, and higher order $\beta_{40}$ deformation, have been taken into account in the evaluated process. Significant improvements can be achieved in training and validating data sets. Especially, based on the $P^{*}$ model, the rms deviations for the training and validation stage fall into $\sigma^{\mathrm{T}}=0.0084$ fm and $\sigma^{\mathrm{V}}=0.0124$ fm, and $\sigma^{\mathrm{TV}}=0.0092$ fm for entire set. $P^{*}$ model can reproduce the shell quenching phenomena and the rapid increase of charge radii across neutron magic numbers. This suggests that $P^{*}$ model persists the predictive power when the theoretical uncertainties are taken into account. Even for nuclei located far from those contained in the training set, $P^{*}$ model can still make available predictions.
This work presents the potential approach of Bayesian neural network in exploring nuclear charge radii. We should mention that physically motivated features play an indispensable role in extrapolating cases far away from the $\beta$-stability line. Although the Monte Carlo dropout Bayesian neural network provides rather small root-mean-square deviation in simulating the theoretical results, the larger deviation can also be encountered in some local regions. This means the underlying physical mechanism cannot be captured adequately in the training and validation data sets. In addition, the number of training data sets has an influence on describing and predicting the nuclear charge radii. Therefore, more complex input structures should be taken into account properly in the proceeding research.
Acknowledgements
This work was supported by the Open Project of Guangxi Key Laboratory of Nuclear Physics and Nuclear Technology, No. NLK2023-05 and the Central Government Guidance Funds for Local Scientific and Technological Development, China (No. Guike ZY22096024).

### v2

@@ III Results and Discussion
For convenience of discussion, we use $P^{*}$ to denote the results obtained by the MC-dropout BNN model with input neurons $z=\{Z,$ $A$, $I^{2}$, $\delta$, $P^{*}$, $\nu_{n}$, $\nu_{p}$, $\beta_{20}$, $\beta_{40}$, $LI\}$. We also give the results obtained by the MC-dropout BNN model with input neurons $z=\{Z,$ $A$, $I^{2}$, $\delta$, $P$, $\nu_{n}$, $\nu_{p}$, $\beta_{20}$, $\beta_{40}$, $LI\}$, which is denoted by $P$ for comparison. Here, we should be mentioned that the constructed $P^{*}$ model is just used to avoid the distorted cases where the valence neutrons or protons are vanished. In the training set, we use the $815$ experimental data given in Ref. ANGELI201369 . The more recent experimental data PhysRevLett.126.032502 ; LI2021101440 , containing 118 data for nuclei with $Z\geq 20$ and $A\geq 40$, are used to test the predictive power of the constructed model in the validation set. In contrast to Refs. PhysRevC.105.014308 ; DONG2023137726 , the training data sets are reduced owing to the whole data of charge radii along nickel isotopic chain are transformed into the validation data set.
Figure 2: Variations of the RMSDs for the train, validation, and total sets under the framework of D4 PhysRevC.105.014308 , D5 DONG2023137726 , D6 DONG2023137726 , D8, P, and P models.
@@
To further quantify the predictive power, the root-mean-square deviation (RMSD) between the results obtained by the MC-dropout BNN model and the corresponding experimental data are calculated as follows,
$\sigma^{(\mathrm{T},\mathrm{V},\mathrm{TV})}=\sqrt{\frac{1}{N_{(\mathrm{T},\mathrm{V},\mathrm{TV})}}\sum_{i=1}^{N_{(\mathrm{T},\mathrm{V},\mathrm{TV})}}\left(R_{\mathrm{ch},i}^{\mathrm{theo}.}-R_{\mathrm{ch},i}^{\mathrm{exp}.}\right)^{2}}.$ (11)
@@
where $N_{\mathrm{T}}$, $N_{\mathrm{V}}$, and $N_{\mathrm{TV}}$ are the numbers of charge radii contained in the training (T), validation (V), and total (TV) data sets, and the subscript $i$ represents the $i$th nucleus in the given data sets.
To show the quantitative analysis, the influence of the input structure on the determination of the predictive power is discussed as follows. Meanwhile, the results obtained with the input structure $z=\{Z,$ $A$, $I^{2}$, $\delta$, $P^{*}$, $\nu_{n}$, $\nu_{p}$, $LI\}$ are also used to perform a comparison research, which is labeled as the D8 model. In contrast to the D8 model, the deformation parameters $\beta_{20}$ and $\beta_{40}$ are introduced into the input structures of the P model. As shown in Fig. 2, the variations of the RMSDs for the train, validation, and total sets under the framework of the D4 PhysRevC.105.014308 , D5 DONG2023137726 , D6 DONG2023137726 , D8, P and P models are plotted. For the D4 model, the values of the RMSDs for the train, validation, and total sets are the largest with respect to other models. This can be understood easily, because the structure phenomena cannot be adequately captured in this model. The D5 and D6 models give almost equivalent values of the RMSDs for the train and the total sets. In the validation set, the value of the RMSD obtained by the D6 model is less than that obtained by the D5 model due to the marked odd-even staggering of charge radii in the abrupt shape-phase staggering regions.
@@
For the D4, D5, and D6 models, the Casten factor associated with the valence neutrons and protons is considered as input parameter. However, shape deformation has not been taken into account in the validated protocol. In recent study, the input sets $\{Z$, $A$, $\delta$, $\nu_{p}$, $\nu_{n}$, $\beta_{20}\}$ have been used in the training and validation data sets, the RMSDs have been reduced to $\sigma^{\mathrm{T}}=0.0106$ fm and $\sigma^{\mathrm{TV}}=0.0130$ fm Zhang10662945 . In contrast to the D6 model, the D8 model gives the lower RMSDs in the train and total data sets. For the validation data set, the RMSD value obtained by the D8 model is slightly larger than that obtained by the D6 model. As demonstrated in Ref. PhysRevC.88.011301 , the RMSD in charge radii can be reduced if the shape deformation is further incorporated into the calibrated protocol. Here, it should be found that the P model further reduces the values of the RMSDs for the train and total data sets compared to the D8 model. Meanwhile, for the validation set, the value of the RMSD obtained by the D8 model is almost equivalent to the result obtained by the P model.
@@
For the P model, the values of the RMSDs for the train, validation, and total data sets are further reduced with respect to those obtained by the P model, or even the results obtained by the D4, D5, D6, and D8 models. In Table 1, the $\sigma^{\mathrm{T}}$, $\sigma^{\mathrm{V}}$, and $\sigma^{\mathrm{TV}}$ values of the P and P models derived from the Eq. (11) are shown, respectively. In our calculations, the valence neutrons, valence protons, Casten factor, and the shape-phase staggerings in the limited region are simultaneously taken into account during the simulated process. Meanwhile, the shape deformation parameters $\beta_{20}$ and $\beta_{40}$ are also included. The deduced value of $\sigma^{\mathrm{TV}}$ deriving from the $P$ model is equivalent to Ref. Zhang10662945 , but less than the results shown in Refs. PhysRevC.105.014308 ; DONG2023137726 . The $\sigma^{\mathrm{TV}}$ value obtained by the $P^{*}$ model can be further reduced to be $0.0090$ fm. Furthermore, the values of the RMSDs for the total set are sequentially reduced from the D4 model to the P model. This means that more input quantities can reduce the epistemic uncertainty in the machine learning model. Consequently, the input structure has an influence on determining the predictive accuracy. Furthermore, it should be mentioned that the revised Casten factor is instructive to predicting the nuclear charge radii.
@@
Figure 3: Nuclear chart displaying the absolute differences (in units of fm) between the rms charge radii obtained by the $P^{*}$ model and the corresponding experimental data in the entire set. The results are obtained from the training and validation sets, respectively.
Significant improvement has been achieved by the $P^{*}$ model in both the interpreted and the extrapolated predictions of nuclear charge radii compared to the $P$ model. To visualize the details about the training and validation of the MC-dropout BNN approach, as shown in Fig. 3, the difference between the calibrated rms charge radii with the $P^{*}$ model and the corresponding experimental data are depicted in the nuclear chart. From this figure, it can be found that most of the experimental data can be reproduced well. However, for some specific regions, the predicted values are heavily deviated from the experimental data. As is well known, the inverse OES or the weakened OES can be evidently observed around Eu ($Z=63$Alkhazov1990 and At($Z=85$PhysRevC.99.054317 isotopes. In our calibrated protocol, the inverse odd-even staggering behavior has not been considered properly. This seems to lead larger deviations between the predicted rms charge radii and the corresponding experimental data. Meanwhile, larger deviations can also be encountered along manganese isotopes. This means that more underlying mechanisms should be properly covered through the input structures.
@@
Figure 4: Root-mean-square charge radii of calcium (a), nickel (b), tin (c), and lead (d) isotopes obtained by the $P^{*}$ and $P$ models compared to the experimental data ANGELI201369 ; LI2021101440 . The results obtained by the np formula ($R_{\mathrm{np}}$) are also drawn for making a guide eye. The data covered by gray area are not contained in the training set, namely the pure predictions. The neutron magic numbers $N=20$, 28, 50, 82, and 126 are emphasized by the vertical lines.
@@
Shell closure effects are naturally observed in the charge radii throughout the whole nuclear chart ANGELI201369 ; LI2021101440 ; PhysRevLett.128.022502 ; PhysRevLett.129.132501 ; PhysRevLett.131.102501 . This is attributed to the rather small isospin dependence of the spin-orbit interactions PhysRevLett.74.3744 . The relativistic mean field theory Geng:2003pk ; PhysRevC.102.024307 ; PhysRevC.109.064302 and the Fayans energy density functional model PhysRevC.95.064328 ; PhysRevLett.122.192502 can reproduce the shell quenching phenomena in nuclear charge radii well. However, as demonstrated in Ref. PhysRevLett.128.022502 , the Skyrme density functional theory cannot describe the shell closure effects well. This provides a potential signature to identify the predictive power of nuclear models in describing the local variations of the charge radii. In Fig. 4, the rms charge radii of calcium (a), nickel (b), tin (c), and lead (d) isotopes are depicted with both the $P^{*}$ and $P$ models, respectively. As mentioned above, MC-dropout BNN model can provide the quantified uncertainty in the validated process. Therefore the corresponding error bars are also shown in Fig. 4. For calcium isotopes, the $P^{*}$ and $P$ models give the similar trend of changes of charge radii. The slight deviations can be found at the neutron-deficient 34Ca and neutron-rich 54,55Ca isotopes. Meanwhile, charge radii of 39,43Ca isotopes can be reproduced well by the $P^{*}$ model compared to the $P$ model. Likewise, the systematical evolution of charge radii along tin isotopes can also be reproduced well by both the $P^{*}$ and $P$ models as shown in Fig. 4 (c).
@@
In Fig. 4 (b), it is obviously found that the $N=28$ shell closure effect cannot be reproduced well by the $P$ model. This leads to heavily overestimated charge radii of the 54-56,58Ni isotopes. The same scenarios can also be found in the 65-68Ni isotopes, but with a slight deviation. Meanwhile, the charge radius of 70Ni obtained by the $P$ model is slightly underestimated compared to $P^{*}$ model. Toward proton-rich regions, the pronounced deviations can be obviously exhibited between the predictive results obtained by the $P^{*}$ and $P$ models. These larger deviations can also be encountered in the neutron-rich lead isotopes as shown in Fig. 4 (d). Here, it should be mentioned that the rapid increase in charge radii can be reproduced well by the $P^{*}$ model, but the $P$ model fails to follow this trend beyond the neutron number $N=126$. All of these results suggest that the $P^{*}$ model can provide a predictive power to describe the nuclear charge radii. Therefore, an extrapolated simulation with high accuracy is necessary. In Table. 2, charge radii of calcium, nickel, tin, and lead isotopes are shown explicitly. These results are encouraged to provide a theoretical guide for the experimental research.
The input parameters that include the modified Casten factor can describe the systematical evolution of nuclear charge radii well. This can be easily found from the rms deviation of the training and validation data sets. Furthermore, as shown in Fig. 4, the shrunken phenomena in charge radii can be reproduced well by the $P^{*}$ model. Actually, the Casten factor represents the neutron-proton correlation in describing the trend of changes of nuclear charge radii. As demonstrated in Refs. PhysRevC.53.1599 ; MILLER2019360 ; COSYN2021136526 , the neutron-proton correlation has an influence on determining the nuclear charge radii. This can also be found in Refs. PhysRevC.102.024307 ; PhysRevC.109.064302 where the neutron-proton correlation derived from the difference between the neutron and proton Cooper pairs condensation can reproduce the shell quenching of nuclear charge radii as well. All of these suggest that more appropriate interactions should be covered in describing the nuclear charge radii. In particular, the difference of charge radii of mirror-pair nuclei can be used to ascertain the nuclear symmetry energy. Thus, more available charge radii data are required in theoretical study.
@@ IV Summary and conclusions
More input parameter sets physically motivated, such as proton number, mass number, isospin-asymmetry degree, pairing effect, shell closure effect, valence neutrons and protons, Casten factor $P$ or $P^{*}$, local shape-phase staggering phenomena, quadrupole deformation $\beta_{20}$, and higher order $\beta_{40}$ deformation, have been taken into account in the evaluated process. Significant improvements can be achieved in training and validating data sets. Especially, the $P^{*}$ model achieves superior performance, with rms deviations of $\sigma^{\mathrm{T}}=0.0084$ fm (train), $\sigma^{\mathrm{V}}=0.0124$ fm (validation) and $\sigma^{\mathrm{TV}}=0.0090$ fm (total). The $P^{*}$ model can reproduce the shell quenching phenomena and the rapid increase of charge radii across neutron magic numbers. This suggests that the $P^{*}$ model persists the predictive power when theoretical uncertainties are taken into account. For nuclei located far from those contained in the training set, the $P^{*}$ model can still make reliable predictions.
This work presents the potential application of Bayesian neural network in exploring nuclear charge radii. We should mention that more physically motivated features play an indispensable role in extrapolating the charge radii of nuclei far away from the $\beta$-stability line. Although the Monte Carlo dropout Bayesian neural network provides rather small root-mean-square deviation in simulating the theoretical results, the larger deviation can also be encountered in some local regions. This means that the underlying physical mechanism cannot be adequately captured in the training and validation data sets. In addition, the number of training data sets has an influence on describing and predicting the nuclear charge radii. Therefore, more underlying input structures should be taken into account properly in the proceeding investigation.
V Acknowledgements
This work was supported by the Open Project of Guangxi Key Laboratory of Nuclear Physics and Nuclear Technology, No. NLK2023-05, the Central Government Guidance Funds for Local Scientific and Technological Development, China (No. Guike ZY22096024), the Natural Science Foundation of Ningxia Province, China (No. 2024AAC03015), and the Key Laboratory of Beam Technology of Ministry of Education, China (No. BEAM2024G04).